OpenStudy (anonymous):

What is a possible first step to eliminate the variable y in the following system? 4x + 6y = 1 5x – 3y = 8

OpenStudy (anonymous):

Tell me: is there any combination of the second equation that we can add to the first equation that will get rid of y?

OpenStudy (anonymous):

multiply the second equation with 2

OpenStudy (andriod09):

try this method: 4x+6y=1 5x-3y=8 9x+3y=8 then solve for x, then fill in the equation and solve for y.